OpenStudy (angela210793):

May some1 help me?Please (1/sqra+1) + (a/sqra+a)

OpenStudy (angela210793):

transform it please....

OpenStudy (amistre64):

this one?

OpenStudy (angela210793):

Yes please :)

OpenStudy (amistre64):

\[\frac{1}{\sqrt{a}+1} + \frac{a}{\sqrt{a}+a}\] is this teh correct interpretation of it?

OpenStudy (amistre64):

sqra+1 is a little to ambiguous for me to decipher correctly; gonna need some guidance ;)

OpenStudy (angela210793):

hmmm....guidance? idk.....the eq. is right how u have written it....:)

OpenStudy (amistre64):

good, then I can procede

OpenStudy (angela210793):

yes u can :) meanwhile can i move from the pc?

OpenStudy (amistre64):

lol.... maybe :)

OpenStudy (angela210793):

good thanks :):) I'll pass by later :) Thanks a lot :):)

OpenStudy (amistre64):

we can 'cross' multiply to get a common denom that works out sqrt(a) + a + asqrt(a) + a ---------------------- (sqrt(a)+1) (sqrt(a)+a) sqrt(a) + a sqrt(a) + 2a ---------------------- a +a +sqrt(a) + a sqrt(a) sqrt(a) + a sqrt(a) + 2a ---------------------- = 1 if I did it right lol 2a +sqrt(a) + a sqrt(a)

OpenStudy (angela210793):

Thanks a loooooooot :)